OpenStudy (anonymous):

FAN AND MEDAL PLEASE HELP Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 4x2 + 6x + 2 = 0

OpenStudy (anonymous):

legggoo raiders haha

OpenStudy (anonymous):

except we didn't make the playoffs :(

OpenStudy (skullpatrol):

http://i.stack.imgur.com/WSqVW.gif

OpenStudy (anonymous):

So true

OpenStudy (anonymous):

can you help please

OpenStudy (anonymous):

It says it wants numerical answers

OpenStudy (anonymous):

answer really

OpenStudy (skullpatrol):

factor out the greatest monomial factor first

OpenStudy (anonymous):

ok lets see

OpenStudy (anonymous):

out of the 4x2 + 6x + 2 = 0?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4 (x2 +1.5x + .5) + 0

OpenStudy (anonymous):

sorry i don't really know how to do this

OpenStudy (skullpatrol):

$$\Huge 4x^2 + 6x + 2 = 0$$$$\Huge 2*2x^2 + 2*3x + 2 = 0$$

OpenStudy (anonymous):

ok i see

OpenStudy (skullpatrol):

What is the greatest common factor?

OpenStudy (anonymous):

2

OpenStudy (skullpatrol):

Ok, now factor it out please.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm sorry can you show me how

OpenStudy (skullpatrol):

$$\Huge ab+ac=a(b+c)$$

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

does a=2

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

b=3?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

2*3 + 2*2 = 2(3+2)

OpenStudy (skullpatrol):